OpenStudy (anonymous):

A ball rolls off a horizontal table 1 m high with a velocity of 3 m/s. Find: a.) The time before the ball strikes the floor b.) The horizontal distance of the ball from the foot of the table

OpenStudy (anonymous):

the direction of initial velocity is x

OpenStudy (anonymous):

so the vertical movement is free fall with the initial velocity 0m/s equation:Y: \[1/2\times g \times t ^{2} =1m\]

OpenStudy (anonymous):

Rearrange haijon's expression to find the time of flight. and then use x=vt with v=3 and the time you've previously solved for to find the horizontal distance.

OpenStudy (anonymous):

t=\[\sqrt{2h/g}\] h=1m g=9.8 m/\[s ^{2}\]